# A causal viewpoint on prediction model performance under changes in case-mix: discrimination and calibration respond differently for prognosis and diagnosis predictions

### v2

@@ 4.2 Empirical Study
As an empirical validation we re-used data from a published systematic review on prediction models in cardiovascular disease which included 2030 external validations of 1382 predictions models [wesslerExternalValidationsCardiovascular2021] and whose data is publicly available at https://www.pacecpmregistry.org. The review investigated changes in model performance when comparing the original publication with later external validation studies. The authors classified the prediction models as either ‘diagnostic’ or ‘prognostic’ (indicated by a follow-up time of less than 3 months, 3-6 months or more than 6 months). Selecting only prediction models with one or more validations and information on AUC in both the original study and validation study, and with information on model type (diagnostic versus prognostic), 1170 validation studies remained of 342 prediction models, 16 of which were validation studies of 11 diagnostic models. Comparing the AUC in the original study ($\text{AUC}_{0}$) with external validation studies ($\text{AUC}_{1}$), we calculated the relative difference in AUC as suggested by the authors:
@@
Our framework predicts that for diagnostic models that predict in the anti-causal direction, the AUC remains the same so $\text{AUC}_{0}=\text{AUC}_{1}$, thus $\text{VAR}(\delta)=0$, but not for prognosis models that predict in the causal direction. The studies in this systematic review are likely not perfectly causal or anti-causal, and because of varying finite sample sizes, variation in AUC will occur. Still we expect the variance of $\delta$ to be higher for prognosis models than for diagnosis models. In these data, using an F-test to compare the to variances this was indeed the case, with $\text{VAR}(\delta_{\text{diagnostic}})=0.019\approx 0.122*\text{VAR}(\delta_{\text{prognostic}})$, p-value$<0.001$. Unfortunately the review provided no quantitative measures of calibration so a similar comparison of the variance of changes in calibration could not be made.
@@ 5 Discussion
Prior work noted that prediction models that are calibrated in multiple environments are provably free from anti-causal predictors [waldCalibrationOutofDomainGeneralization2021]. Our focus is in the reverse direction: when to expect stable calibration across environments. Bareinboim and Pearl describe the data fusion problem and discuss when a causal effect may be transportable across settings and with what data [bareinboimCausalInferenceDatafusion2016]. Our work describes when certain specific functionals from the distributions are stable across settings, tailored to typical needs in the (medical) prediction model setting.
Our framework also provides a new perspective on the results of the study by Fehr at al [fehrAssessingTransportabilityClinical2023]. They experimented with prediction models that contained either causal factors of the outcome (related to our prognosis models), anti-causal factors (related to our diagnosis models), or a combination of both. The performance of different prediction models was evaluated under different shifts in variables that were at the same time a direct cause of the outcome and a cause of other variables. Fehr et al find that for models predicting only with cause variables, the calibration is stable under interventions on only cause variables, as directly explained by our framework. When predicting with anti-causal factors, they observed that under interventions on the cause variables, the calibration degrades for models that are well calibrated on the training data. This setting is the closest to our diagnostic setting, though technically it is a mix of the anti-causal DAG 2(b) and confounding 2(c).
@@
Limitations are that the definition of a shift in case-mix is an abstraction and pure interventions on only either the features or the outcome may be unrealistic in practice. Many diagnostic prediction models may contain features that have a causal path to the diagnosis (e.g. age), or ‘risk factors’ for the disease that are not caused by the presence or absence of the disease. Systematic reviews of diagnostic models indeed show variation in sensitivity and specificity with variation in disease prevalence, a phenomenon also referred to as the spectrum-effect [leeflangVariationTestSensitivity2013]. Still, when compared with prognostic models, diagnostic models had lower variability in discrimination in our empirical study. The current empirical validation was limited, and classifying diagnostic models as anti-causal and prognostic models as causal may be too crude. Also, no quantitative data on calibration were available to test whether calibration was more stable for prognostic models. Future empirical studies of externally evaluated prediction models will shed more light on how this theory pans out in practice. Mis-calibration may occur when variables not included in the model are also shifted between environments.
Still, the framework is useful as it explains prior observations and provides guidance to evaluators of prediction models on what to expect and how to explain certain changes changes in prediction model performance.
@@ Acknowledgments
The author kindly acknowledges Anne de Hond, Oisín Ryan and Valentijn de Jong for fruitful discussions on this framework.
@@
Let $P(X,Y)$ be a joint distribution over feature $X$ and binary outcome $Y$, and $f:X\to[0,1]$ a prediction model. $f$ is perfectly calibrated with respect to $P(X,Y)$ if, for all $\alpha\in[0,1]$ in the range of $f$, $\operatorname{\mathbb{E}}_{X,Y\sim P(X,Y)}[Y|f(X)=\alpha]=\alpha$.
@@
Given prediction model $f:\mathcal{X}\to[0,1]$ and environment $E\in\{\text{train},\text{test}\}$. Assume $P_{E}(Y,X)=P_{E}(X)P(Y|X)$ (shift in $P(X)$ but not $P(Y|X)$ between environments). Furthermore assume $\text{supp}(P_{\text{test}}(X))\subseteq\text{supp}(P_{\text{train}}(X))$ (marginal distribution of $X$ in training data is non-zero wherever the test distribution is non-zero). Define the miscalibration of $f$ under $P_{E}$ for value $X=x$ as:
$\xi(x):=|f(x)-P(Y|X=x)|$ (1)
Then the integrated calibration index (ICI) on distribution $P_{E}$ is [austinIntegratedCalibrationIndex2019]:
@@
Theorem statement: a perfectly calibrated model (i.e. $\xi(x)=0,\forall x\in\text{supp}(P_{\text{train}}(X))$) remains perfectly calibrated under marginal shifts in $X$, i.e.
@@
$=\int_{\text{supp}{P_{\text{test}}(X)}}\phi_{\text{test}}(x)\xi(x)$ (5)
$=\int_{\text{supp}{P_{\text{test}}(X)}}\phi_{\text{test}}(x)*0$ (6)
$=0\int_{\text{supp}{P_{\text{test}}(X)}}\phi_{\text{test}}(x)$ (7)
@@
Given prediction model $f:\mathcal{X}\to[0,1]$ and environment $E\in\{\text{train},\text{test}\}$. Assume $P_{E}(Y,X)=P_{E}(Y)P(X|Y)$ (marginal shift of $Y$ but not $X|Y$). Furthermore assume $10<P_{E}(Y=1)<$ 1 (marginal distribution of $Y$ in both distributions is non-deterministic). Then for all thresholds $\tau$:
$sens_{\text{test}}(\tau)$ $=sens_{\text{train}}(\tau)$ (10)
$spec_{\text{test}}(\tau)$ $=spec_{\text{train}}(\tau)$ (11)
@@
$AUC_{\text{train}}=AUC_{\text{train}}$ (12)

### v3

@@ 4.2 Empirical Study
As an empirical evaluation we re-used data from a published systematic review on prediction models in cardiovascular disease which included 2030 external validations of 1382 predictions models [wesslerExternalValidationsCardiovascular2021] and whose data is publicly available at https://www.pacecpmregistry.org. The review investigated changes in model performance when comparing the original publication with later external validation studies. The authors classified the prediction models as either ‘diagnostic’ or ‘prognostic’ (indicated by a follow-up time of less than 3 months, 3–6 months or more than 6 months). Selecting only prediction models with one or more validations and information on AUC in both the original study and validation study, and with information on model type (diagnostic versus prognostic), 1170 validation studies remained of 342 prediction models, 16 of which were validation studies of 11 diagnostic models. Comparing the AUC in the original study ($\text{AUC}_{0}$) with external validation studies ($\text{AUC}_{1}$), we calculated the relative difference in AUC as suggested by the authors:
@@
Our framework predicts that for diagnostic models that predict in the anti-causal direction, the AUC remains the same so $\text{AUC}_{0}=\text{AUC}_{1}$, thus $\text{VAR}(\delta)=0$, but not for prognosis models that predict in the causal direction. The studies in this systematic review are likely not perfectly causal or anti-causal, and because of sampling variance, variation in AUC will occur. Still we expect the variance of $\delta$ to be higher for prognosis models than for diagnosis models. In these data this was indeed the case with $\text{VAR}(\delta_{\text{prognostic}})\approx 8.2*\text{VAR}(\delta_{\text{diagnostic}})=0.019$, 95% confidence of ratio: 3.41 - 15.10, p-value for F-test $<0.001$. Unfortunately the review provided no quantitative measures of calibration so a similar comparison of the variance of changes in calibration could not be made.
The code needed to reproduce the simulation experiment and empirical evaluation are available at Zenodo.
5 Related work
The notion that calibration is stable under shifts in the distribution of the cause variables has long been appreciated (e.g. [schoelkopfCausalAnticausalLearning2012, piccininniDirectedAcyclicGraphs2020, waldCalibrationOutofDomainGeneralization2021]). Much of the theory in this paper is inspired by Schölkopf’s work on causal and anti-causal learning [schoelkopfCausalAnticausalLearning2012]. This work connects the general framework to the medical setting in two ways: we define a familiar term ‘case-mix’ in a formal causal language, and then derive how two canonical metrics of predictive performance (discrimination and calibration) respond differently to changes in case-mix for causal prediction models (prognosis) and anti-causal models (diagnosis).
Prior work noted that prediction models that are calibrated in multiple environments are provably free from anti-causal predictors [waldCalibrationOutofDomainGeneralization2021]. Our focus is in the reverse direction: when to expect stable calibration across environments. Jaladoust and colleagues derived general bounds for functionals of the target distribution in a new environment [jalaldoustPartialTransportabilityDomain2024]. Our work describes when certain specific functionals from the distributions (discrimination and calibration) are stable across settings, tailored to typical needs in the (medical) prediction model setting. Other prior work requires detailed assumptions on the causal relationships between variables [subbaswamyUnifyingCausalFramework2022], or access to data from multiple environments. Subbaswamy consideres loss functions of the form $l(\hat{y},y)$ which implicitly depend on the joint distribution of $X,Y$ trough the expectation $L=E_{X,Y}[l(\hat{y},y)]$. Our work is focused on metrics that are direct functionals of the conditional distribution $Y|X$ and $X|Y$. Also, Subbaswamy focusses on min-max optimality.
Our framework also provides a new perspective on the results of the study by Fehr at al [fehrAssessingTransportabilityClinical2023]. They experimented with prediction models that contained either causal factors of the outcome (related to our prognosis models), anti-causal factors (related to our diagnosis models), or a combination of both. The performance of different prediction models was evaluated under different shifts in variables that were at the same time a direct cause of the outcome and a cause of other variables. Fehr et al found that for models predicting only with cause variables, the calibration is stable under interventions on only cause variables, as directly explained by our main result. When predicting with anti-causal factors, they observed that under interventions on the cause variables, the calibration degrades for models that are well calibrated on the training data. This setting is the closest to our diagnostic setting, though technically it is a mix of the anti-causal DAG 2(b) and the confounded DAG 2(c).
@@ 6 Discussion
Limitations are that the definition of a shift in case-mix is an abstraction and pure interventions on only either the features or the outcome may be unrealistic in practice. Many diagnostic prediction models may contain features that have a causal path to the diagnosis (e.g. age), or ‘risk factors’ for the disease that are not caused by the presence or absence of the disease. Also, the ‘disease’ itself may be an abstraction, and the diagnosis used in medical practice may be a combination of effects of an underlying biological process. Systematic reviews of diagnostic models indeed show variation in sensitivity and specificity with variation in disease prevalence, a phenomenon also referred to as the spectrum-effect [leeflangVariationTestsSensitivity2013]. Still, when compared with prognostic models, diagnostic models had lower variability in discrimination in our empirical study. The current empirical evaluation was limited, and classifying diagnostic models as anti-causal and prognostic models as causal may be too crude. Also, no quantitative data on calibration were available to test whether calibration was more stable for prognostic models. Future empirical studies of externally evaluated prediction models will shed more light on how this theory pans out in practice. Mis-calibration may occur when variables not included in the model are also shifted between environments.
What are the implications of the causal case-mix framework for different stakeholders? A main use of this new causal case-mix framework is to provide an explanation of (lack of) expected and observed differences in prediction model performance across environments. For prediction model developers, this framework provides a new way to think about the features included in a prediction model. In some settings such as triaging patients in emergency room for early medical evaluation, the utility of a prediction model depends mostly on its discrimination. In other settings such as cardiovascular risk management, a prediction model’s utility depends on its calibration. Depending on this utility function, prediction model developers may opt to include mostly causal or anti-causal features in a prediction model, if dependable performance across environments is desired. The framework adds another perspective on the discussion on when and where to re-calibrate a prediction model [swaminathanReflexiveRecalibrationCausal2025, lekadirFUTUREAIInternationalConsensus2025]. When calibration is important and the model does include anti-causal features, it is likely necessary to always recalibrate the model when taking it to a new environment, in line with many recommendations [youssefExternalValidationAI2023]. However, when discrimination changes for a prediction model in the causal direction, this may not warrent a re-fitting of the model as this change is to be expected under a shift in case-mix. A decrease (or increase) of discrimination may be indicative of a shift in the data, meaning more (or less) patients with probabilities closer to 50%, than of a bad model.
For researchers that evaluate prediction models and policy makers, it was long known that no models are robust to arbitrary changes in distribution. This framework implies that a subset of models should have stable calibration or discrimination. For example, observing a stable discrimination of a diagnostic model with anti-causal features in several different environments may provide confidence that the model is indeed robust to environmental changes. At the same time, this model’s calibration should not be stable across evaluations. Whereas normally a stable calibration would be seen as a re-assuring sign, having both calibration and discrimination stable across environments is a sign that the environments are not meaningfully different at all. A stable discrimination paired with unstable calibration (or the other way around) is a stronger sign of robustness to changes in environment than when both are stable, as this would only occur when the environment are too similar.
@@ Acknowledgments
The author kindly acknowledges Anne de Hond, Oisin Ryan and Valentijn de Jong for fruitful discussions on this framework.
@@
Let $P(X,Y)$ be a joint distribution over feature $X$ and binary outcome $Y$, and $f:\mathcal{X}\to[0,1]$ a deterministic prediction model. $f$ is perfectly calibrated with respect to $P(X,Y)$ if, for all $\alpha\in[0,1]$ in the range of $f$, $\operatorname{\mathbb{E}}_{X,Y\sim P(X,Y)}[Y|f(X)=\alpha]=\alpha$.
@@
Given binary $Y$, prediction model $f:\mathcal{X}\to[0,1]$ and environment $E\in\{\text{train},\text{test}\}$. Assume $X$ takes on values from a measureable space $\mathcal{X}$ with measures $\phi_{train}(x),\phi_{test}(x)$ on the training and testing environment, and assume $\phi_{test}(x)<<\phi_{train}(x)$ (the support of the test distribution is contained in the support of the training distribution). Assume $P_{E}(Y,X)=P_{E}(X)P(Y|X)$ (shift in $P(X)$ but not $P(Y|X)$ between environments). Define the miscalibration of $f$ under $P_{E}$ for value $X=x$ as:
$\xi(x):=|f(x)-P(Y=1|X=x)|$ (1)
Then the integrated calibration index (ICI) [austinIntegratedCalibrationIndex2019] on distribution $P_{E}$ is:
@@
Theorem statement: a model that is perfectly calibrated on the training distribution (i.e. $\xi(x)=0\iff\phi_{train}(x)>0$) remains perfectly calibrated in the test distribution:
@@
By assumption we have $\phi_{train}(x)>0\implies\xi(x)=0$, because $\phi_{train}(x)>0\implies\phi_{test}(x)>0$ we also have that $\phi_{test}(x)>0\implies\xi(x)=0$ ($f$ is calibrated for all values of $x$ in the test distribution). Denote $\text{supp}_{test}(X)$ the subset of $\mathcal{X}$ where $\phi_{test}(x)>0$. By definition of ICI we have that
@@
[…]
@@
Given binary outcome $Y$, prediction model $f:\mathcal{X}\to[0,1]$ and environment $E\in\{\text{train},\text{test}\}$. Assume $P_{E}(Y,X)=P_{E}(Y)P(X|Y)$ (marginal shift of $Y$ but not $X|Y$). Furthermore assume $10<P_{E}(Y=1)<$ 1 (marginal distribution of $Y$ in both distributions is non-deterministic). Then for all thresholds $0\leq\tau\leq 1$:
$\text{sens}_{\text{test}}(\tau)$ $=\text{sens}_{\text{train}}(\tau)$ (10)
$\text{spec}_{\text{test}}(\tau)$ $=\text{spec}_{\text{train}}(\tau)$ (11)
@@
$AUC_{\text{train}}=AUC_{\text{test}}$ (12)